40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 577     Entry time: Tue Apr 12 19:49:39 2011
Author: tara 
Type: DailyProgress 
Category: BEAT 
Subject: Beat measurement is back 

I realigned the beam on the beat path and measured the beat. With new eddy current damping, the peaks at 3.73 Hz now becomes smaller.

beat_2011_04_12.png

 

ELOG V3.1.3-